# A Survey on Graph Neural Networks for Time Series: Forecasting, Classification, Imputation, and Anomaly Detection

### v1

@@ Abstract
Time series are the primary data type used to record dynamic system measurements and generated in great volume by both physical sensors and online processes (virtual sensors). Time series analytics is therefore crucial to unlocking the wealth of information implicit in available data. With the recent advancements in graph neural networks (GNNs), there has been a surge in GNN-based approaches for time series analysis. Approaches can explicitly model inter-temporal and inter-variable relationships, which traditional and other deep neural network-based methods struggle to do. In this survey, we provide a comprehensive review of graph neural networks for time series analysis (GNN4TS), encompassing four fundamental dimensions: Forecasting, classification, anomaly detection, and imputation. Our aim is to guide designers and practitioners to understand, build applications, and advance research of GNN4TS. At first, we provide a comprehensive task-oriented taxonomy of GNN4TS. Then, we present and discuss representative research works and, finally, discuss mainstream applications of GNN4TS. A comprehensive discussion of potential future research directions completes the survey. This survey, for the first time, brings together a vast array of knowledge on GNN-based time series research, highlighting both the foundations, practical applications, and opportunities of graph neural networks for time series analysis.
@@ 1 Introduction
The advent of advanced sensing and data stream processing technologies has led to an explosion of time series data, one of the most ubiquitous data types that captures and records activity across a wide range of domains [1, 2, 3]. The analysis of time series data not only provides insights into past trends but also facilitates a multitude of tasks such as forecasting [4], classification [5], anomaly detection [6], and data imputation [7]. This lays the groundwork for time series modeling paradigms that leverage on historical data to understand current and future possibilities. Time series analytics have become increasingly crucial in various fields, including but not limited to cloud computing, transportation, energy, finance, social networks, and the Internet-of-Things [8, 9, 10, 11].
@@
Figure 1: Graph neural networks for time series analysis (GNN4TS). In this example of wind farm, different analytical tasks can be categorized into time series forecasting, classification, anomaly detection, and imputation.
Many time series involve complex interactions across time (such as lags in propagation of effects) and variables (such as the relationship among the variables representing neighboring traffic sensors). By treating time points or variables as nodes and their relationships as edges, a model structured in the manner of a network or graph can effectively learn the intricacies of these relationships. Indeed, much time series data is spatial-temporal in nature, with different variables in the series capturing information about different locations – space – too, meaning it encapsulates not only time information but also spatial relationships [12]. This is particularly evident in scenarios such as urban traffic networks, population migration, and global weather forecasting. In these instances, a localized change, such as a traffic accident at an intersection, an epidemic outbreak in a suburb, or extreme weather in a specific area, can propagate and influence neighboring regions. This might manifest as increased traffic volume on adjacent roads, the spread of disease to neighboring suburbs, or altered weather conditions in nearby areas. This spatial-temporal characteristic is a common feature of many dynamic systems, including another example of the wind farm in Figure 1, where the underlying time series data displays a range of correlations and heterogeneities [13]. These factors contribute to the formation of complex and intricate patterns, posing significant challenges for effective modeling. Traditional analytic tools, such as support vector regression (SVR) [14, 15], gradient boosting decision tree (GBDT) [16, 17], vector autoregressive (VAR) [18, 19], and autoregressive integrated moving average (ARIMA) [20, 21], struggle to handle complex time series relations (e.g., nonlinearities and inter-series relationships), resulting in less accurate prediction results [22]. The advent of deep learning technologies in the past decade has led to the development of different neural networks based on convolutional neural networks (CNN) [23, 24], recurrent neural networks (RNN) [25, 26], and Transformers [27], which have shown significant advantages in modeling real-world time series data. However, one of the biggest limitations of the above methods is that they do not explicitly model the spatial relations existing between time series in non-Euclidean space [13, 28], which limits their expressiveness [28].
In recent years, graph neural networks (GNNs) have emerged as a powerful tool for learning non-Euclidean data representations [29, 30, 31], paving the way for modeling real-world time series data. This enables the capture of diverse and intricate relationships, both inter-variable (connections between different variables within a multivariate series) and inter-temporal (dependencies between different points in time). Considering the complex spatial-temporal dependencies inherent in real-world scenarios, a line of studies has integrated GNNs with various temporal modeling frameworks to capture both spatial and temporal dynamics and demonstrate promising results [13, 32, 33, 34, 35]. This modeling approach has also been widely adopted in many real-world application sectors with different time series data, including transportation [36], on-demand services [37, 38], energy [39], healthcare [40, 41], economy [42], and other fields [43, 44, 45]. While early research efforts were primarily concentrated on various forecasting scenarios [13, 33, 34], recent advancements in time series analysis utilizing GNNs have demonstrated promising outcomes in other mainstream tasks. These include classification [46, 47], anomaly detection [48, 49], and imputation [50, 51]. In Figure 1, we provide an overview of graph neural networks for time series analysis (GNN4TS).
Related Surveys. Despite the growing body of research related to performing various time series analytic tasks with GNNs, existing surveys are relatively limited in number and tend to focus on specific perspectives within a restricted scope. For instance, the survey by Wang et al. [12] offers a review of deep learning techniques for spatial-temporal data mining, but it does not specifically concentrate on GNN-based methods and fails to reflect the most recent advancements in the field. The survey by Ye et al. [35] zeroes in on graph-based deep learning architectures in the traffic domain, primarily considering different forecasting scenarios. A recent survey by Jin et al.[13] offers an overview of GNNs for predictive learning in urban computing, but neither extends its coverage to other application domains nor thoroughly discusses other tasks related to time series analysis. Finally, we mention the work by Rahmani et al. [36], which expands the survey of GNNs to many intelligent transportation systems, but tasks other than forecasting remain overlooked. A detailed comparison between our survey and others is presented in Table I.
@@
Survey Domain Scope
@@
Wang et al. [12]
[…]
Ye et al. [35]
[…]
Jiang and Luo [33]
[…]
Bui et al. [34]
[…]
Jin et al. [13]
[…]
Al Sahili and Awad [32]
[…]
Rahmani et al. [36]
[…]
Our Survey
@@
To fill the gap, this survey offers a comprehensive and up-to-date review of graph neural networks for time series analysis, encompassing mainstream tasks ranging from time series forecasting, classification, anomaly detection, and imputation. Specifically, we first provide two broad views to classify and discuss existing works from the task- and methodology-oriented perspectives. Then, we delve into six popular application sectors within the existing research of GNN4TS, and propose several potential future research directions. Our survey is intended for general machine learning practitioners interested in exploring and keeping abreast of the latest advancements in graph neural networks for time series analysis. It is also suitable for domain experts seeking to apply GNN4TS to new applications or explore novel possibilities building on recent advancements. The key contributions of our survey are summarized as follows:
@@
  • Unified and Structured Taxonomy. We present a unified framework to structurally categorize existing works from task- and methodology-oriented perspectives. In the first classification, we provide an overview of tasks in time series analysis, including various forecasting, classification, anomaly detection, and imputation settings that are targeted in most GNN-based related work. We further present a structured taxonomy in the second classification to dissect the graph neural networks for time series analysis from the perspective of spatial and temporal dependencies modeling, as well as the overall model architecture.
@@
  • Insight into Future Research Directions. We provide an overview of potential future research directions, offering insights and suggestions that could guide and inspire future work in the field of GNN4TS.
The remainder of this survey is organized as follows: section 2 provides important notations and related definitions used throughout the paper. section 3 presents the taxonomy of GNN4TS from different perspectives, along with a general pipeline. section 4, section 5, section 6, and section 7 review four mainstream analytic tasks in the GNN4TS literature. section 8 surveys popular applications of GNN4TS across various fields, while section 9 examines open questions and potential future directions. Finally, section 10 concludes this survey.
@@ 2 Definition and Notation
In this section, we provide the definitions and notations used throughout the paper. In the subsequent text, we employ bold uppercase letters (e.g., $\mathbf{X}$), bold lowercase letters (e.g., $\mathbf{x}$), and calligraphic letters (e.g., $\mathcal{V}$) to denote matrices, vectors, and sets, respectively.
In this survey, we start by defining time series data, which serves as a fundamental basis for abstracting various real-world systems, such as photovoltaic and traffic networks. Time series data comprises a sequence of observations gathered or recorded over a period of time. This data can be either regularly or irregularly sampled, with the latter also referred to as time series data with missing values. Within each of these cases, the data can be further classified into two primary types: univariate and multivariate time series.
@@
A univariate time series is a sequence of scalar observations collected over time, which can be regularly or irregularly sampled. A regularly-sampled univariate time series is defined as $\mathbf{X}=\{x_{1},x_{2},...,x_{T}\}\in\mathbb{R}^{T}$, where $x_{t}\in\mathbb{R}$ for $t=1,2,\dots,T$. For an irregularly-sampled univariate time series, the observations are collected at non-uniform time intervals, such as $\mathbf{X}=\{(t_{1},x_{1}),(t_{2},x_{2}),...,(t_{T},x_{T})\}\in\mathbb{R}^{T}$, where time points are non-uniformly spaced.
@@
A multivariate time series is a sequence of vector observations collected over time, i.e., $\mathbf{X}\in\mathbb{R}^{N\times T}$, where each vector comprises $N$ variables. Similarly, it can either be regularly or irregularly sampled. A regularly-sampled multivariate time series has vector observations collected at uniform time intervals, i.e., $\mathbf{x}_{t}\in\mathbb{R}^{N}$. In an irregularly-sampled multivariate time series, there may be $N$ unaligned time series with respect to time steps, which means that there are $0\leq n\leq N$ observations available at each time step.
The majority of research based on GNNs focuses on modeling multivariate time series, as they can be naturally abstracted into spatial-temporal graphs. This abstraction allows for an accurate characterization of dynamic inter-temporal and inter-variable dependencies. The former describes the relations between different time steps within each time series (e.g., the temporal dynamics of red nodes between $t_{1}$ and $t_{3}$ in Figure 2), while the latter captures dependencies between time series (e.g., the spatial relations between four nodes at each time step in Figure 2), such as the geographical information of the sensors generating the data for each variable. We first define static attributed graphs as follows.
@@
An attributed graph is a graph that associates each node with a set of attributes, representing node features. Formally, an attributed graph is defined as $\mathcal{G}=(\mathbf{A},\mathbf{X})$, which consists of a (weighted) adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ and a node-feature matrix $\mathbf{X}\in\mathbb{R}^{N\times D}$. The adjacency matrix represents the graph topology, which can be characterized by node and edge sets $\mathcal{V}$ and $\mathcal{E}$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of $N$ nodes, and =$\mathcal{E}=\{e_{ij}:=(v_{i},v_{j})\in\mathcal{V}\times\mathcal{V}\mid\mathbf{A}_{ij}\neq 0\}$ is the set of edges, with $\mathbf{A}_{ij}$ being the $(i,j)$-th entry in the adjacency matrix $\mathbf{A}$. The feature matrix $\mathbf{X}$ contains the node attributes, where the $i$-th row $\mathbf{x}_{i}\in\mathbb{R}^{D}$ represents the $D$-dimensional feature vector of node $v_{i}$.
@@
A spatial-temporal graph can also be interpreted as discrete-time dynamic graphs [52]. Formally, we define it as $\mathcal{G}=\{\mathcal{G}_{1},\mathcal{G}_{2},\cdots,\mathcal{G}_{T}\}$, where $\mathcal{G}_{t}=(\mathbf{A}_{t},\mathbf{X}_{t})$ for each time step $t$. $\mathbf{A}_{t}\in\mathbb{R}^{N\times N}$ is the adjacency matrix representing the graph topology at time $t$, and $\mathbf{X}_{t}\in\mathbb{R}^{N\times D}$ is the feature matrix containing the node attributes at time $t$. The adjacency matrix may either evolve over time or remain fixed, depending on specific settings. When abstracting time series data, we let $\mathbf{X}_{t}:=\mathbf{x}_{t}\in\mathbb{R}^{N}$.
Figure 2: Examples of spatial-temporal graphs.
We introduce graph neural networks as modern deep learning models to process graph-structured data. The core operation in typical GNNs, often referred to as graph convolution, involves exchanging information across neighboring nodes. In the context of time series analysis, this operation enables us to explicitly rely on the time series dependencies represented by the graph edges. Aware of the different nuances, we define GNNs in the spatial domain, which involves transforming the input signal with learnable functions along the dimension of $N$.
@@
We adopt the definition presented in [30]. Given an attributed graph $\mathcal{G}=(\mathbf{A},\mathbf{X})$, we define $\mathbf{x}_{i}=\mathbf{X}[i,:]\in\mathbb{R}^{D}$ as the $D$-dimensional feature vector of node $v_{i}$. A GNN learns node representations through two primary functions: $\textsc{Aggregate}(\cdot)$ and $\textsc{Combine}(\cdot)$. The $\textsc{Aggregate}(\cdot)$ function computes and aggregates messages from neighboring nodes, while the $\textsc{Combine}(\cdot)$ function merges the aggregated and previous self-information to transform node embeddings. Formally, the $k$-th layer in a GNN is defined by convolution
@@
The above formulation in Equation 1 is referred to as spatial GNNs, as opposed to spectral GNNs which defines convolution from the lens of spectral graph theory. We refer the reader to recent publication [28] for a deeper analysis of spectral versus spatial GNNs, and [29, 53] for a comprehensive review of GNNs.
@@
Heuristic-based graph. This group of methods extracts graph structures from data based on various heuristics, such as:
@@
  • Pairwise Connectivity: In this approach, the graph structure is determined by the connectivity between pairs of nodes, like that determined by transportation networks where the adjacency matrix $\mathbf{A}$ is defined as:
@@
  • Functional Dependence: This approach defines the graph structure based on the functional dependence between pairs of nodes. These include the construction of adjacency matrix $\mathbf{A}$ based on Granger causality [58]:
@@
    Other examples involve transfer entropy (TE) [59] and directed phase lag index (DPLI) [60].
Learning-based graph. In contrast to heuristic-based methods, learning-based approaches aim to directly learn the graph structure from the data end-to-end with the downstream task. These techniques typically involve optimizing the graph structure alongside model parameters during the training process, e.g., embedding-based [61], attention-based [62, 63], sampling-based [64, 65]. These learning-based approaches enable the discovery of more complex and potentially more informative graph structures compared to heuristic-based methods.

### v3

@@ Abstract
Time series are the primary data type used to record dynamic system measurements and generated in great volume by both physical sensors and online processes (virtual sensors). Time series analytics is therefore crucial to unlocking the wealth of information implicit in available data. With the recent advancements in graph neural networks (GNNs), there has been a surge in GNN-based approaches for time series analysis. These approaches can explicitly model inter-temporal and inter-variable relationships, which traditional and other deep neural network-based methods struggle to do. In this survey, we provide a comprehensive review of graph neural networks for time series analysis (GNN4TS), encompassing four fundamental dimensions: forecasting, classification, anomaly detection, and imputation. Our aim is to guide designers and practitioners to understand, build applications, and advance research of GNN4TS. At first, we provide a comprehensive task-oriented taxonomy of GNN4TS. Then, we present and discuss representative research works and introduce mainstream applications of GNN4TS. A comprehensive discussion of potential future research directions completes the survey. This survey, for the first time, brings together a vast array of knowledge on GNN-based time series research, highlighting foundations, practical applications, and opportunities of graph neural networks for time series analysis.
@@
Project Page: https://github.com/KimMeen/Awesome-GNN4TS
@@ 1 Introduction
The advent of advanced sensing and data stream processing technologies has led to an explosion of time series data [1, 2, 3]. The analysis of time series not only provides insights into past trends but also facilitates a multitude of tasks such as forecasting [4], classification [5], anomaly detection [6], and data imputation [7]. This lays the groundwork for time series modeling paradigms that leverage on historical data to understand current and future possibilities. Time series analytics have become increasingly crucial in various fields, including but not limited to cloud computing, transportation, energy, finance, social networks, and the Internet-of-Things [8, 9, 10].
@@
Figure 1: Graph neural networks for time series analysis (GNN4TS). In this example of a wind farm, different analytical tasks can be categorized into time series forecasting, classification, anomaly detection, and imputation.
@@
Survey Year Domain Scope
@@
[…]
    Specifically,  represents “Not Covered”,  signifies “Partially Covered”, and  corresponds to “Fully Covered”.
Many time series involve complex interactions across time (such as lags in propagation of effects) and variables (such as the relationship among the variables representing neighboring traffic sensors). By treating time points or variables as nodes and their relationships as edges, a model structured in the manner of a network or graph can effectively solve the task at hand by exploiting both data and relational information. Indeed, much time series data is spatial-temporal in nature, with different variables in the series capturing information about different locations – space – too, meaning it encapsulates not only time information but also spatial relationships [11]. This is particularly evident in scenarios such as urban traffic networks, population migration, and global weather forecasting. In these instances, a localized change, such as a traffic accident at an intersection, an epidemic outbreak in a suburb, or extreme weather in a specific area, can propagate and influence neighboring regions. This spatial-temporal characteristic is a common feature of many dynamic systems, including the wind farm in Figure 1, where the underlying time series displays a range of correlations and heterogeneities [15, 18]. Traditional analytic tools, such as support vector regression (SVR) [19], gradient boosting decision tree (GBDT) [20], vector autoregressive (VAR) [21], and autoregressive integrated moving average (ARIMA) [22], struggle to handle complex time series relations (e.g., nonlinearities and inter-variable relationships), resulting in less accurate prediction results [23]. The advent of deep learning technologies has led to the development of different neural networks based on convolutional neural networks (CNN) [24, 25], recurrent neural networks (RNN) [26], and transformers [27], which have shown significant advantages in modeling real-world time series data. However, one of the biggest limitations of the above methods is that they do not explicitly model the spatial relations existing between time series in non-Euclidean space [15], which limits their expressiveness [28].
In recent years, graph neural networks (GNNs) have emerged as a powerful tool for learning non-Euclidean data representations [29, 30, 31, 32], paving the way for modeling real-world time series data. This enables the capture of diverse and intricate relationships, both inter-variable (connections between different variables within a multivariate series) and inter-temporal (dependencies between different points in time). Considering the complex spatial-temporal dependencies inherent in real-world scenarios, a line of studies has integrated GNNs with various temporal modeling frameworks to capture both spatial and temporal dynamics and demonstrate promising results [15, 16, 13, 14, 12]. While early research efforts were primarily concentrated on various forecasting scenarios [15, 13, 14], recent advancements in time series analysis utilizing GNNs have demonstrated promising outcomes in other mainstream tasks. These include classification [33, 34], anomaly detection [35, 36], and imputation [37, 38]. In Figure 1, we provide an overview of graph neural networks for time series analysis (GNN4TS).
Related Surveys. Despite the growing body of research performing various time series analytic tasks with GNNs, existing surveys tend to focus on specific perspectives within a restricted scope. For instance, the survey by Wang et al. [11] offers a review of deep learning techniques for spatial-temporal data mining, but it does not specifically concentrate on GNN-based methods. The survey by Ye et al. [12] zeroes in on graph-based deep learning architectures in the traffic domain, primarily considering forecasting scenarios. A recent survey by Jin et al.[15] offers an overview of GNNs for predictive learning in urban computing, but neither extends its coverage to other application domains nor thoroughly discusses other tasks related to time series analysis. Finally, we mention the work by Rahmani et al. [17], which expands the survey of GNNs to many intelligent transportation systems, but tasks other than forecasting remain overlooked. A detailed comparison between our survey and others is presented in section 1.
To fill the gap, this survey offers a comprehensive and up-to-date review of graph neural networks for time series analysis, encompassing the majority of tasks ranging from time series forecasting, classification, anomaly detection, and imputation. Specifically, we first provide two broad views to classify and discuss existing works from the task- and methodology-oriented perspectives. Then, we delve into six popular application sectors within the existing research of GNN4TS, and propose several potential future research directions. The key contributions of our survey are summarized as follows:
@@
  • Unified and Structured Taxonomy. We present a unified framework to structurally categorize existing works from task- and methodology-oriented perspectives. In the first classification, we offer an overview of tasks in time series analysis, covering different problem settings prevalent in GNN-based research; and in the second classification, we dissect GNN4TS in terms of spatial and temporal dependencies modeling and the overall model architecture.
@@
  • Future Research Directions. We shed light on potential future research directions, offering insights and suggestions that could guide and inspire future research in the field of GNN4TS.
The remainder of this survey is organized as follows: section 2 provides notations used throughout the paper. section 3 presents the taxonomy of GNN4TS from different perspectives. section 4, section 5, section 6, and section 7 review the four major tasks in the GNN4TS literature. section 8 surveys popular applications of GNN4TS across various fields, while section 9 examines open questions and potential future directions.
@@ 2 Definition and Notation
Time series data comprises a sequence of observations gathered or recorded over a period of time. This data can be either regularly or irregularly sampled, with the latter also referred to as time series data with missing values. Within each of these cases, the data can be further classified into two primary types: univariate and multivariate time series. In the sequel, we employ bold uppercase letters (e.g., $\mathbf{X}$), bold lowercase letters (e.g., $\mathbf{x}$), and calligraphic letters (e.g., $\mathcal{V}$) to denote matrices, vectors, and sets, respectively.
@@
A univariate time series is a sequence of scalar observations collected over time, which can be regularly or irregularly sampled. A regularly sampled univariate time series is defined as $\mathbf{X}=\{x_{1},x_{2},...,x_{T}\}\in\mathbb{R}^{T}$, where $x_{t}\in\mathbb{R}$. For an irregularly sampled univariate time series, observations are collected at non-uniform time intervals, such as $\mathbf{X}=\{(t_{1},x_{1}),(t_{2},x_{2}),...,(t_{T},x_{T})\}\in\mathbb{R}^{T}$, where time points are non-uniformly spaced.
@@
A multivariate time series is a sequence of $N$-dimensional vector observations collected over time, i.e., $\mathbf{X}\in\mathbb{R}^{N\times T}$. A regularly sampled multivariate time series has vector observations collected at uniform time intervals, i.e., $\mathbf{x}_{t}\in\mathbb{R}^{N}$. In an irregularly sampled multivariate time series, there are possibly $N$ unaligned time series with respect to time steps, which implies only $0\leq n\leq N$ observations available at each time step.
The majority of research based on GNNs focuses on modeling multivariate time series, as they can be naturally abstracted into spatial-temporal graphs. This abstraction allows for an accurate characterization of dynamic inter-temporal and inter-variable dependencies. The former describes the relations between different time steps within each time series (e.g., the temporal dynamics of red nodes between $t_{1}$ and $t_{3}$ in Figure 2), while the latter captures dependencies between time series (e.g., the spatial relations between four nodes at each time step in Figure 2), such as the geographical information of the sensors generating the data for each variable. To illustrate this, we first define attributed graphs.
@@
An attributed graph is a static graph that associates each node with a set of attributes, representing node features. Formally, an attributed graph is defined as $\mathcal{G}=(\mathbf{A},\mathbf{X})$, which consists of a (weighted) adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ and a node-feature matrix $\mathbf{X}\in\mathbb{R}^{N\times D}$. The adjacency matrix represents the graph topology, which can be characterized by $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$, the set of $N$ nodes, and =$\mathcal{E}=\{e_{ij}:=(v_{i},v_{j})\in\mathcal{V}\times\mathcal{V}\mid\mathbf{A}_{ij}\neq 0\}$, the set of edges; $\mathbf{A}_{ij}$ is the $(i,j)$-th entry in the adjacency matrix $\mathbf{A}$. The feature matrix $\mathbf{X}$ contains the node attributes, where the $i$-th row $\mathbf{x}_{i}\in\mathbb{R}^{D}$ represents the $D$-dimensional feature vector of node $v_{i}$.
@@
A spatial-temporal graph can be interpreted as a discrete-time dynamic graph [39, 31], i.e., $\mathcal{G}=\{\mathcal{G}_{1},\mathcal{G}_{2},\cdots,\mathcal{G}_{T}\}$, where $\mathcal{G}_{t}=(\mathbf{A}_{t},\mathbf{X}_{t})$ denotes an attributed graph at time $t$. $\mathbf{A}_{t}\in\mathbb{R}^{N\times N}$ and $\mathbf{X}_{t}\in\mathbb{R}^{N\times D}$ are corresponding adjacency and feature matrices. $\mathbf{A}_{t}$ may either evolve over time or remain fixed, depending on specific settings. When abstracting time series data, we let $\mathbf{X}_{t}:=\mathbf{x}_{t}\in\mathbb{R}^{N}$.
Figure 2: Examples of spatial-temporal graphs, where node colors represent distinct features. The top and bottom panels demonstrate spatio-temporal graphs with fixed and dynamic graph structures over time, respectively.
We introduce graph neural networks as modern deep learning models to process graph-structured data. The core operation in typical GNNs, often referred to as graph convolution, involves exchanging information across neighboring nodes. In the context of time series analysis, this operation enables us to explicitly rely on the inter-variable dependencies represented by the graph edges. Aware of the different nuances, we define GNNs in the spatial domain, which involves transforming the input signal with learnable functions along the dimension of $N$.
@@
Given an attributed graph $\mathcal{G}=(\mathbf{A},\mathbf{X})$, we define $\mathbf{x}_{i}=\mathbf{X}[i,:]\in\mathbb{R}^{D}$ as the $D$-dimensional feature vector of node $v_{i}$. A GNN learns node representations through two primary functions [40]: $\textsc{Aggregate}(\cdot)$ and $\textsc{Combine}(\cdot)$. The $\textsc{Aggregate}(\cdot)$ function computes and aggregates messages from neighboring nodes, while the $\textsc{Combine}(\cdot)$ function merges the aggregated and previous states to transform node embeddings. Formally, the $k$-th layer in a GNN is defined by the extended
@@
The above formulation in Equation 1 is referred to as spatial GNNs, as opposed to spectral GNNs which defines convolution from the lens of spectral graph theory. We refer the reader to recent publication [28] for a deeper analysis of spectral versus spatial GNNs, and [29] for a comprehensive review of GNNs.
@@
Heuristic-based Graphs. This group of methods extracts graph structures from data based on heuristics, such as:
@@
  • Pairwise Connectivity: In this approach, the graph structure is determined by the connectivity between pairs of nodes, like that determined by transportation networks. The adjacency matrix $\mathbf{A}$ is defined as:
@@
  • Functional Dependence: This approach defines the graph structure based on known functional dependencies between pairs of nodes, such as direct causal relationships or dependence from common hidden factors. For instance, adjacency matrix $\mathbf{A}$ can be constructed based on Granger causality [45] as
@@
    Other examples involve transfer entropy (TE) [46] and directed phase lag index (DPLI) [47]. While overlap with previous heuristics exists, functional relations typically represent or estimate actual node dependencies in the data-generating process.
Learning-based Graphs. In contrast to heuristic-based methods, learning-based approaches aim to learn the graph structure directly from the data and end-to-end with the downstream task. Accordingly, graph $\mathbf{A}$ can be defined as a function $\rho({}\cdot{})$ of some trainable model parameters $\Theta$ and, possibly, also of the time series observations $\mathbf{X}$, i.e., $\mathbf{A}=\rho(\Theta,\mathbf{X})$. Embedding-based approaches (e.g., [48, 49]) define the presence of each edge by comparing learned embedding vectors of the associated nodes; as an example, consider $\mathbf{A}_{i,j}=\text{ReLU}(\Theta_{i}^{\top}\Theta_{j})$ with $\Theta_{i},\Theta_{j}$ node embedding of nodes $i,j$, respectively. Another common example involves defining $\mathbf{A}$ through attention scores computed between node signals [50]. Sparsification methods – such as the ReLU activation above or top-k selection – are often applied to discard edges with null or low weights, thereby reducing the computational load requested by dense GNN operations [51]. An alternative strategy is modeling $\mathbf{A}$ as a discrete random variable whose parametric distribution, say $p_{\Theta}(\mathbf{A}\mid\mathbf{X})$, is learned alongside the other model parameters [52, 53]. Learning-based approaches enable the data-driven discovery of less obvious graph structures that are tailored to solve the given task, potentially providing more effective relations than heuristic-based graphs.